Two angles are supplementary. The measure of one angle is one-third the measure of the other. Find the measures of both angles.
Answer by stanbon(75887)

Two angles are supplementary. The measure of one angle is one-third the measure of the other. Find the measures of both angles.
-----
Equations:
x + y = 180
x = (1/3)y
----
Subtritute for "x" and solve for "y"::
(1/3)y + y = 180
(4/3)y = 180
(1/3)y = 45
y = 135 degrees
x = 180-135 = 45 degrees
-------------
Cheers,
Stan H.
--------------